# Efficient Segmentation of Search Markets

### v2

@@ Abstract
I consider a two-sided frictional search market where buyers search and match to vertically differentiated sellers. The market is segmented into submarkets based on seller types, with segmentation serving as a public signal that directs buyers’ search. I characterize the socially efficient and equilibrium allocations of buyers across submarkets for any fixed market segmentation, and identify a Hosios condition under which the equilibrium allocation is efficient. I further examine the design of surplus-maximizing segmentations, demonstrating the role of search externalities in determining whether the constrained-efficient segmentation fully reveals seller types or pools types into at most a binary partition. These results clarify the conditions under which the provision of public information is welfare enhancing in search markets with externalities.
@@
I would like to express my deepest appreciation to Eddie Dekel, Federico Echenique, Bruno Strulovici, and Asher Wolinsky for their guidance and mentorship. This work has also greatly benefited from conversations with Alessandro Bonatti, Odilon Câmara, Hector Chade, Yeon-Koo Che, Pedro Dal Bó, Laura Doval, Jack Fanning, Kyungmin Kim, Stelios Michalopoulos, Bobak Pakzad-Hurson, João Ramos, Jesse Shapiro, Nikhil Vellodi, and Rajiv Vohra. This paper supersedes a previous draft entitled “Random versus Directed Search for Scarce Resources.”
@@ 1.  Introduction
The first set of results characterizes both the socially efficient and equilibrium allocations of buyers across submarkets for any fixed market segmentation. From a welfare perspective, tighter submarkets are advantageous because they facilitate more meetings, thereby increasing the overall volume of trade. However, with a finite measure of buyers, increasing the tightness of one submarket comes at the expense of others. The unique socially efficient allocation resolves this trade-off by equalizing the marginal value of an additional meeting across all submarkets. In other words, a socially efficient allocation balances the thick market externality—where adding a buyer to a submarket increases overall trade opportunities—against the expected surplus generated from trade within that submarket.
In contrast, the unique equilibrium allocation arises from how buyers trade off each submarket’s congestion against the share of expected surplus they capture in that submarket conditional on trade. Specifically, buyers follow a cutoff strategy based on a reservation value: they avoid submarkets where the expected payoff falls below this value. Among the remaining submarkets, those offering higher payoffs conditional on trade attract more buyers until their level of congestion precisely offsets their potential gains from trade.
@@ 1.1.  Related Literature
This paper focuses on the former, and specifically, considers the role of ex-ante information in a market for search goods—goods whose quality is revealed upon inspection. [^2] In this paper, a buyer observes a seller’s type after the two meet, which reveals the quality of the seller’s product. In this context, Anderson and Renault (2006), Choi et al. (2019), and Lyu (2023) analyze a monopolist’s optimal provision of information when buyers use it to decide between visiting the monopolist or taking their outside option. Similarly, Choi et al. (2018) explores the role of ex-ante information in oligopoly markets for horizontally differentiated search goods. However, these studies focus on settings with either a single buyer or sellers without capacity constraints, thereby abstracting away from search externalities. This paper addresses this gap by focusing on how ex-ante information interacts with thick market externalities and congestion externalities in markets with multiple buyers and capacity-constrained sellers. [^3] Of course, other types of externalities could be considered. For example. Vellodi (2018) considers the role of ex-ante information in the entry and exit decisions of sellers in a search market, and shows that suppressing information can incentivize the entry of new sellers and delay the exit of incumbents.
@@
This work also connects to the extensive literature on competitive search markets, where sellers post prices that buyers use to direct their search. [^4] While the labor and matching literature often uses the terms directed search and competitive search interchangeably, I follow Wright et al. (2021) in distinguishing between the two. As per their definition, “Directed search means agents see some, although perhaps not all, characteristics of other agents, and based on that choose where to look for counterparties,” while “competitive search equilibrium means that agents on one side of the market post the terms of trade, while agents on the other side observe what is posted and direct their search accordingly.” Thus, competitive search is a special case of directed search. Examples include search for exchange goods (Butters, 1977; Peters, 1991), labor markets (Montgomery, 1991; Moen, 1997; Mortensen and Wright, 2002), and markets with two-sided heterogeneity (Shi, 2002; Shimer, 2005; Eeckhout and Kircher, 2010). A central insight from this literature is that equilibrium outcomes are typically efficient because posted prices internalize search externalities.
@@
In contrast, this paper considers directed search driven solely by information rather than posted prices. Terms of trade are instead determined via bargaining once a buyer and seller meet, [^5] One such environment is the labor market for high-skilled workers: firms often direct their search by recruiting candidates from top universities, but salaries are negotiated after a successful completion of the interview process. Similarly, the posted-price framework is unsuitable in environments where sellers lack commitment power, or where prices are fixed exogenously. as in Diamond (1982), Mortensen (1982), and Pissarides (1985). Consequently, prices cannot fully internalize the search externalities unless a Hosios condition (Hosios, 1990) is satisfied. Hence, this paper demonstrates that beyond price-based competitive search, directed search can give rise to inefficient equilibrium outcomes. Furthermore, it identifies conditions under which information design can mitigate these inefficiencies, offering insights into the optimal provision of ex-ante information in search markets.
@@
Few papers explore directed search beyond price-based competitive search. The most relevant work in this regard is Menzio (2007), which examines a private-values labor market where informed firms send cheap-talk messages that uninformed workers use to direct their search. Once a trading partner is met, the terms of trade are determined via bargaining. Menzio (2007) shows that while informative cheap-talk equilibria exist, a fully revealing equilibrium does not, and all equilibrium outcomes are inefficient. [^6] When the trading mechanism is a first-price auction rather than bargaining, Kim and Kircher (2015) demonstrates the existence of a fully revealing and efficient cheap-talk equilibrium in a private-values setting, whereas Kim (2012) shows that no such equilibrium exists in a setting with vertical differentiation. These inefficiencies stem from two interdependent externalities: search externalities on the workers’ side of the market, and signaling externalities on the firms’ side, which arise because the informational content of a firm’s cheap-talk message depends on the messaging strategies of other firms.
@@
In this paper, however, the market segmentation (and thus, its informational content) is taken as exogenous for both buyers and sellers, reflecting the increasingly data-driven segmentation of markets based on certification or reviews rather than cheap talk. As a result, signaling externalities are absent from my model. Unlike Menzio (2007), isolating the interaction between information and search externalities demonstrates that fully revealing seller types is not inherently incompatible with efficient equilibria; both coexist under a Hosios-like condition.
@@
Finally, this paper loosely relates to the body of work on the role of interim information, especially in markets for experience goods—goods whose quality is revealed only upon consumption. In the context of large search markets, Lauermann (2012) and Lester et al. (2019) demonstrate the negative welfare implications of interim information provision. A growing body of work also considers the design of interim information for experience goods in single-agent settings: Romanyuk and Smolin (2019), Dogan and Hu (2022), Hu (2022), Mekonnen et al. (2023), and Sato and Shirakawa (2023) all consider centralized provision of interim information, while Board and Lu (2018), Au and Whitmeyer (2023), Mekonnen and Pakzad-Hurson (2024), and He and Li (2023) consider decentralized provision of information. Beyond search markets, the welfare implications and the design of interim information has also been analyzed in various other contexts, including price-discriminating monopoly (Bergemann et al., 2015), differentiated Bertrand competition (Elliott et al., 2021), bilateral matching markets Condorelli and Szentes (2023), and many-to-one matching markets Li (2023).
@@ 2.  Model
A two-sided market is populated by a unit mass of homogeneous buyers on one side and a mass $k>0$ of heterogeneous sellers on the other. Each buyer (he) has a unit demand for a good, and each seller (she) has a unit supply. A seller’s type, denoted by $\theta$, captures the quality of her good. The type space $\Theta$ is a compact interval normalized to $[0,1]$. Types are distributed according to an absolutely continuous probability measure $\mu\in\Delta(\Theta)$. [^7] Given a set $X$, $\mathcal{B}(X)$ denotes the Borel $\sigma$-algebra of $X$, and $\Delta(X)$ denotes the space of probability measures over $(X,\mathcal{B}(X))$. Each seller knows her own type, but each buyer ex-ante knows only the type distribution; he observes a seller’s type only upon meeting her.
The market on the sellers’ side is segmented into submarkets, with the segmentation taken as exogenous and common knowledge among the buyers and sellers. Each seller is allocated to a submarket based on her type, possibly with some randomization. Each buyer, on the other hand, can freely choose which submarkets to join.
@@ 2.1.  Market segmentation
A market segmentation is given by a stochastic kernel $\sigma:\mathcal{B}(\Delta(\Theta))\times\Theta\to[0,1]$, where for each $\theta\in\Theta$, the mapping $B\mapsto\sigma(B,\theta)$ is a probability on $(\Delta(\Theta),\mathcal{B}(\Delta(\Theta))$; for each $B\in\mathcal{B}(\Delta(\Theta))$, the mapping $\theta\mapsto\sigma(B,\theta)$ is $\mathcal{B}(\Theta)$-measurable; and for all $A\in\mathcal{B}(\Theta)$ and $B\in\mathcal{B}(\Delta(\Theta))$,
$\int_{B\times\Theta}\nu(A)\sigma(d\nu,\theta)\mu(d\theta)=\int_{A}\sigma(B,\theta)\mu(d\theta).$
The last condition imposes a Bayes-rule structure. Intuitively, each $\nu\in\Delta(\Theta)$ can be thought of as a “label” for a submarket. Each type-$\theta$ seller is allocated to a submarket labeled “$\nu$” with probability $\sigma(d\nu,\theta)$. The equality in (2.1) implies that the type distribution within a submarket labeled “$\nu$” is in fact $\nu\in\Delta(\Theta)$, and that this type distribution is derived by Bayes’ rule. For example, a perfect segmentation, denoted $\sigma^{PS}$, corresponds to a market segmentation where each seller type is allocated to its own unique submarket, i.e., $\sigma^{PS}(\cdot,\theta)=\delta_{\delta_{\theta}}(\cdot)$ for each $\theta\in\Theta$ [^8] Given some space $X$, the Dirac measure $\delta_{x}$ satisfies $\delta_{x}(B)=1$ whenever $x\in B\in\mathcal{B}(X)$.
Under this interpretation, submarkets serve a dual role for buyers: each submarket is a platform for meeting a seller, and it also serves as a signal of the seller types therein. This is perhaps clearer to see by defining $\sigma_{\mu}(d\nu)\coloneqq\int_{\Theta}\sigma(d\nu,\theta)\mu(d\theta)$ and evaluating (2.1) at $B=\Delta(\Theta)$, which yields
@@ 2.2.  Bilateral Meeting functions
Formally, a submarket populated by a measure $b$ of buyers and a measure $s$ of sellers has a submarket tightness of $t\coloneqq b/s\in\mathbb{R}_{+}\cup\{\infty\}$. Each seller meets a buyer with probability $m(t)$ and meets no one with probability $1-m(t)$, where the mapping $t\mapsto m(t)$ denotes the meeting function. As a measure $s\cdot m(t)$ of sellers meet buyers and meetings are bilateral, there must also be a measure $s\cdot m(t)$ of buyers that meet sellers. Thus, each buyer meets a seller with probability $m(t)/t$. The following assumptions on the meeting function are maintained for the remainder of the paper.
@@
The first assumption implies that both $m(t)$ and $m(t)/t$ are well-defined probabilities. To motivate the second assumption, notice that as the submarket tightness increases, the market contains even more buyers relative to sellers, which has two effects: On the one hand, each seller in the submarket has more opportunities to meet a buyer, giving rise to a thick market externality on the sellers’ side. This effect is captured by the monotonicity of the meeting function. On the other hand, each buyer in the submarket faces more competition to meet a seller, giving rise to a congestion externality on the buyers’ side. This effect is captured by the strict concavity of the meeting function, which guarantees that the mapping $t\mapsto m(t)/t$ is strictly decreasing.
@@
While Assumption 1 is common in the search and matching literature, it is admittedly restrictive. However, by standard continuity arguments, the results of this paper extend to meeting functions that do not satisfy Assumption 1, provided such functions can be approximated by meeting functions that do satisfy it. For example, consider the meeting function $m(t)=\min\{\alpha,\beta t\}$, which generalizes the frictionless meeting function given by $m(t)=\min\{1,t\}$. This meeting function does not satisfy the second point of Assumption 1, as it is only weakly increasing, weakly concave, and non-differentiable at $t=\alpha/\beta$. Nevertheless, the CES meeting function approximates the frictionless one as $\rho\to\infty$, and thus, the insights of this paper extend to this class of generalized frictionless meeting functions.

### v3

# How to Segment a Search Market: Information Design and Directed Search

@@ Abstract
This paper examines when the public provision of information in search markets improves welfare. I consider a two-sided frictional search market in which buyers match with vertically differentiated sellers. The market is segmented into submarkets based on seller types. Such segmentation serves as a public signal that buyers use to direct their search. Given a segmentation, I characterize both the socially efficient and the equilibrium allocation of buyers across submarkets, and identify a Hosios-type condition under which the equilibrium allocation is efficient. I then analyze the design of surplus-maximizing segmentations, showing that the nature of search externalities determines when the constrained-efficient segmentation fully separates seller types or pools them into at most a binary partition.
@@
Keywords: directed search, market segmentation, information design I would like to express my deepest appreciation to Eddie Dekel, Federico Echenique, Bruno Strulovici, and Asher Wolinsky for their guidance and mentorship. This work has also greatly benefited from conversations with Alessandro Bonatti, Odilon Câmara, Hector Chade, Yeon-Koo Che, Pedro Dal Bó, Laura Doval, Jack Fanning, Kyungmin Kim, Stelios Michalopoulos, Bobak Pakzad-Hurson, João Ramos, Jesse Shapiro, Nikhil Vellodi, and Rajiv Vohra. Finally, I thank the editor and anonymous referees for comments and suggestions that greatly improved the paper. This paper supersedes a previous draft entitled “Random versus Directed Search for Scarce Resources.”
@@ 1.  Introduction
The first set of results characterizes both the socially efficient and equilibrium allocations of buyers across submarkets for any fixed market segmentation. From a welfare perspective, tighter submarkets are advantageous because they facilitate more meetings, thereby increasing the overall volume of trade. However, with a finite measure of buyers, raising tightness in one submarket necessarily reduces it elsewhere. The unique socially efficient allocation therefore balances the thick market externality created by tighter submarkets against the expected surplus generated within each submarket.
In contrast, the unique equilibrium allocation arises from how buyers trade off the congestion in each submarket against the share of surplus they expect to capture in that submarket conditional on trade. In equilibrium, buyers follow a cutoff strategy based on a reservation value. They avoid submarkets where the expected payoff falls below this threshold. Among the remaining submarkets, those offering higher payoffs conditional on trade attract more buyers until their level of congestion precisely offsets their potential gains from trade.
@@ 1.1. Related Literature
This paper focuses on the former, and specifically, considers the role of ex-ante information in a market for search goods—goods whose quality is revealed upon inspection. [^2] In this paper, a buyer observes a seller’s type after the two meet, which reveals the quality of the seller’s product. In this context, Anderson and Renault (2006), Choi et al. (2019), and Lyu (2023) analyze a monopolist’s optimal provision of information when buyers use it to decide between visiting the monopolist or taking their outside option. Similarly, Choi et al. (2018) explores the role of ex-ante information in oligopoly markets for horizontally differentiated search goods. However, these studies focus on settings with either a single buyer or sellers without capacity constraints, thereby abstracting away from search externalities. This paper addresses this gap by focusing on how ex-ante information interacts with thick market and congestion externalities in markets with multiple buyers and capacity-constrained sellers. [^3] Of course, other types of externalities could be considered. For example. Vellodi (2018) considers the role of ex-ante information in the entry and exit decisions of sellers in a search market, and shows that suppressing information can incentivize the entry of new sellers and delay the exit of incumbents.
@@
This work also connects to the extensive literature on competitive search markets, where sellers post prices that buyers use to direct their search. [^4] While the labor and matching literature often uses the terms directed search and competitive search interchangeably, I follow Wright et al. (2021) in distinguishing between the two. As per their definition, “Directed search means agents see some, although perhaps not all, characteristics of other agents, and based on that choose where to look for counterparties,” while “competitive search equilibrium means that agents on one side of the market post the terms of trade, while agents on the other side observe what is posted and direct their search accordingly.” Thus, competitive search is a special case of directed search. Examples include search for exchange goods (Butters, 1977; Peters, 1991), labor markets (Montgomery, 1991; Moen, 1997; Mortensen and Wright, 2002), and markets with two-sided heterogeneity (Shi, 2002; Shimer, 2005; Eeckhout and Kircher, 2010a). A central insight from this literature is that equilibrium outcomes are typically efficient because posted prices internalize search externalities.
@@
In contrast, this paper considers directed search driven solely by information rather than posted prices. Terms of trade are instead determined via bargaining once a buyer and seller meet, [^5] One such environment is the labor market for high-skilled workers: firms often direct their search by recruiting candidates from top universities, but salaries are negotiated after a successful completion of the interview process. Similarly, the posted-price framework is unsuitable in environments where sellers lack commitment power, or where prices are fixed exogenously. as in Diamond (1982a), Mortensen (1982), and Pissarides (1985). Consequently, prices cannot fully internalize the search externalities unless a Hosios condition (Hosios, 1990) is satisfied. Hence, this paper demonstrates that beyond price-based competitive search, directed search can give rise to inefficient equilibrium outcomes. Furthermore, it identifies conditions under which information design can mitigate these inefficiencies, offering insights into the optimal provision of ex-ante information in search markets.
@@
Few papers explore directed search beyond price-based competitive search. The most relevant work in this regard is Menzio (2007), who examines a private-values labor market where informed firms send cheap-talk messages that uninformed workers use to direct their search. Once a trading partner is met, the terms of trade are determined via bargaining. Menzio (2007) shows that while informative cheap-talk equilibria exist, all equilibrium outcomes are inefficient. [^6] When the trading mechanism is a first-price auction rather than bargaining, Kim and Kircher (2015) demonstrates the existence of a fully revealing and efficient cheap-talk equilibrium in a private-values setting, whereas Kim (2012) shows that no such equilibrium exists in a setting with vertical differentiation. These inefficiencies stem from two interdependent externalities: search externalities on the workers’ side of the market, and signaling externalities on the firms’ side, which arise because the informational content of a firm’s cheap-talk message depends on the messaging strategies of other firms.
@@
In this paper, however, the market segmentation (and thus, its informational content) is taken as exogenously given by both buyers and sellers, reflecting the increasingly data-driven segmentation of markets based on certification or reviews rather than cheap talk. As a result, signaling externalities are absent from my model, leading to a clean analysis of how information interacts with search externalities. This, in turn, allows me to identify conditions under which efficient and equilibrium outcomes coincide.
Of course, directed search need not be based solely on posted prices or information; rather, it may be based on both. Eeckhout and Kircher (2010b) show that when buyers and sellers meet bilaterally, it is efficient to perfectly segment the market by type, and that this outcome can be decentralized by having buyers direct their search based on both the market segmentation and sellers’ posted prices. Similarly, Cai et al. (2017) demonstrate that bilateral meetings are both sufficient and necessary for perfect market segmentations to emerge as socially efficient. In this respect, my finding that perfect segmentation is efficient is not surprising. However, I show that this outcome can be implemented as an equilibrium if and only if the Hosios condition identified in the paper is satisfied.
@@
Finally, this paper loosely relates to the body of work on the role of interim information, especially in markets for experience goods—goods whose quality is revealed only upon consumption. In the context of large search markets, Lauermann (2012) and Lester et al. (2019) demonstrate the negative welfare implications of interim information provision. A growing body of work also considers the design of interim information for experience goods in single-agent settings: Romanyuk and Smolin (2019), Dogan and Hu (2022), Hu (2022), Mekonnen et al. (2023), and Sato and Shirakawa (2023) all consider centralized provision of interim information, while Board and Lu (2018), Au and Whitmeyer (2023), Mekonnen and Pakzad-Hurson (2024), and He and Li (2023) consider competitive provision of information. Beyond search markets, the welfare implications and the design of interim information has also been analyzed in various other contexts, including price-discriminating monopoly (Bergemann et al., 2015), differentiated Bertrand competition (Elliott et al., 2021), bilateral matching markets Condorelli and Szentes (2023), and many-to-one matching markets Li (2023).
@@ 2.  Model
A two-sided market is populated by a unit mass of homogeneous buyers on one side and a mass $k>0$ of heterogeneous sellers on the other. Each buyer (he) has a unit demand for a good, and each seller (she) has a unit supply. A seller’s type, denoted by $\theta$, captures the quality of her good. The type space $\Theta$ is a compact interval normalized to $[0,1]$. Let $\mathcal{B}(\Theta)$ denote the Borel $\sigma$-algebra of $\Theta$ and $\Delta(\Theta)$ denote the space of probability measures over $(\Theta,\mathcal{B}(\Theta))$. Types are distributed according to an absolutely continuous probability measure $\mu\in\Delta(\Theta)$. Each seller knows her own type, but each buyer ex-ante knows only the type distribution; he observes a seller’s type only upon meeting her.
The market is segmented into submarkets, with the segmentation taken as exogenous and common knowledge among the buyers and sellers. Each seller is allocated to a submarket based on her type, possibly with some randomization. Each buyer, on the other hand, can freely choose which submarkets to join.
@@ 2.1. Market segmentation
A submarket is simply a market populated by a subset of seller types. Accordingly, each submarket can be indexed by the distribution of seller types it contains; that is, a submarket with seller-type distribution $\nu\in\Delta(\Theta)$ is referred to as submarket $\nu$.
A market segmentation is an allocation of sellers to submarkets based on their types. Formally, let $\mathcal{B}(\Delta(\Theta))$ denote the Borel $\sigma$-algebra over $\Delta(\Theta)$, and let $\Delta(\Delta(\Theta))$ be the space of probability measures over $(\Delta(\Theta),\mathcal{B}(\Delta(\Theta)))$. A market segmentation is defined as a stochastic kernel $\sigma:\mathcal{B}(\Delta(\Theta))\times\Theta\to[0,1]$ such that,
  1. $(a)$
    For each $\theta\in\Theta$, the mapping $B\mapsto\sigma(B,\theta)$ is a probability measure in $\Delta(\Delta(\Theta))$;
  2. $(b)$
    For each $B\in\mathcal{B}(\Delta(\Theta))$, the mapping $\theta\mapsto\sigma(B,\theta)$ is $\mathcal{B}(\Theta)$-measurable.
Under segmentation $\sigma$, a type-$\theta$ seller is allocated to submarket $\nu$ with probability $\sigma(d\nu,\theta)$. Consequently, out of the total mass $k$ of sellers, the fraction allocated to submarket $\nu$ is given by
$\sigma_{\mu}(d\nu)\coloneqq\int_{\Theta}\sigma(d\nu,\theta)\mu(d\theta),$
and the support $\operatorname*{supp}(\sigma_{\mu})$ represents the submarkets present in the market.
Of course, the market segmentation must ensure that each submarket $\nu$ actually contains seller types distributed according to $\nu$. In particular, $\sigma$ must satisfy the following consistency condition: for all $A\in\mathcal{B}(\Theta)$ and $B\in\mathcal{B}(\Delta(\Theta))$,
$\int_{A}\sigma(B,\theta)\mu(d\theta)=\int_{B}\nu(A)\sigma_{\mu}(d\nu).$
This identity has a natural interpretation: consider sellers types in some measurable subset $A\subseteq\Theta$. The fraction of these types allocated to submarkets in $B$, given by the left-hand side, must equal the fraction of the same types contained in submarkets in $B$, given by the right-hand side.
The consistency condition in (2.1) guarantees that each submarket $\nu$ indeed has seller-type distribution $\nu$, and that buyers can infer this using Bayes’ rule. Thus, submarkets serve a dual role for buyers: each one is a platform for meeting a seller, and it also serves as a signal of the seller types therein. This is especially transparent when evaluating the consistency condition (2.1) at $B=\Delta(\Theta)$, which yields
@@ 2.2. Bilateral Meeting functions
Formally, a submarket populated by a measure $b$ of buyers and a measure $s$ of sellers has a submarket tightness of $t\coloneqq b/s\in\mathbb{R}_{+}\cup\{\infty\}$. Each seller meets a buyer with probability $m(t)$ and meets no one with probability 1m(t)1-m(t), where the mapping $t\mapsto m(t)$ denotes the meeting function. As a measure $s\cdot m(t)$ of sellers meet buyers and meetings are bilateral, there must also be a measure $s\cdot m(t)$ of buyers that meet sellers. Thus, each buyer meets a seller with probability $m(t)/t$. The following assumptions on the meeting function are maintained for the remainder of the paper.
@@ 2. $()$
The first assumption implies that both $m(t)$ and $m(t)/t$ are well-defined probabilities. To motivate the second assumption, notice that as the submarket tightness increases, the market contains even more buyers relative to sellers, which has two effects: On the one hand, each seller in the submarket has more opportunities to meet a buyer, giving rise to a thick market externality on the sellers’ side. This effect is captured by the monotonicity of the meeting function. [^7] Some earlier work in the matching and search literature uses the term “thick market externality” to refer to increasing returns to scale in the meeting function, which can give rise to multiple equilibrium outcomes in search markets (Diamond, 1982b). In contrast, and consistent with much of the recent literature, I assume constant returns to scale. See Petrongolo and Pissarides (2001) for a more in-depth survey. On the other hand, each buyer in the submarket faces more competition to meet a seller, giving rise to a congestion externality on the buyers’ side. This effect is captured by the strict concavity of the meeting function, which guarantees that the mapping $t\mapsto m(t)/t$ is strictly decreasing.
@@
While Assumption 1 is common in the search and matching literature, it is admittedly restrictive. However, by standard continuity arguments, the results of this paper extend to meeting functions that do not satisfy Assumption 1, provided such functions can be approximated by meeting functions that do satisfy it. For example, consider the meeting function $m(t)=\min\{\alpha,\beta t\}$, which generalizes the frictionless meeting function given by $m(t)=\min\{1,t\}$. This meeting function does not satisfy Point $(b)$ of Assumption 1, as it is only weakly increasing, weakly concave, and non-differentiable at $t=\alpha/\beta$. Nevertheless, the CES meeting function approximates the frictionless one as $\rho\to\infty$, and thus, the insights of this paper extend to this class of generalized frictionless meeting functions.